Project

General

Profile

Actions

MEMO Ops 29112016 » History » Revision 3

« Previous | Revision 3/4 (diff) | Next »
Boyd, Steven, 29 November 2016 13:55


User cycle : 2016/03 - Nov 15th - Dec 16th

  • MOM : Ed Overton who has (today) handed over to Melissa Uchida.
  • MICE is fully staffed (24 hours / day) to the end of the user cycle.
  • Been operating since Nov 17th
    • Nov 18 - Nov 23 : LiH materials programme
    • Nov 24 - Nov 28 : Straight track running
    • Nov 29 - : LiH materials + emittance reduction measurements
  • Allowing BLOCs to be shifters at the same time has helped with BLOC scheduling
  • Running has been largely smooth.
  • Major issues to date :
    • ISIS instability
    • One of the pumps supplying cooling water to the conventional magnets failed leading to continual conventional magnet trips (problem has been understood and mitigation in place, if not fixed)
  • Operational (not to mention personal) concerns over non-delivery of UK grant and imminent disappearance of expert personnel.

Updated by Boyd, Steven almost 7 years ago · 3 revisions